### v2

# Efficient Continual Pre-training by Mitigating the Stability Gap

Abstract
Continual pre-training has increasingly become the predominant approach for adapting Large Language Models (LLMs) to new domains. This process involves updating the pre-trained LLM with a corpus from a new domain, resulting in a shift in the training distribution. To study the behavior of LLMs during this shift, we measured the model’s performance throughout the continual pre-training process. we observed a temporary performance drop at the beginning, followed by a recovery phase, a phenomenon known as the "stability gap," previously noted in vision models classifying new classes. The substantial performance drop and slow recovery associated with this gap lead to inefficient pre-training for domain performance improvement and the forgetting of general task knowledge. To address this issue and enhance LLM performance within a fixed compute budget, we propose three effective strategies: (1) Continually pre-training the LLM on a subset with a proper size for multiple epochs, resulting in faster performance recovery than pre-training the LLM on a large corpus in a single epoch; (2) Pre-training the LLM only on high-quality sub-corpus, which rapidly boosts domain performance; and (3) Using a data mixture similar to the pre-training data to reduce distribution gap. We conduct various experiments on Llama-family models to validate the effectiveness of our strategies in both medical continual pre-training and instruction tuning. For example, our strategies improve the average medical task performance of the OpenLlama-3B model from 36.2% to 40.7% with only 40% of the original training budget and enhance the average general task performance without causing forgetting. Furthermore, we apply our strategies to continually pre-train and instruction-tune the Llama-3-8B model. The resulting model, Llama-3-Physician, achieves the best medical performance among current open-source models, and performs comparably to or even better than GPT-4 on several medical benchmarks. We release our models at https://huggingface.co/YiDuo1999/Llama-3-Physician-8B-Instruct.
1 Introduction
Continual pre-training is an important approach for LLMs to improve their performance in target domains (Huang et al., 2023; Yang et al., 2024a; Chen et al., 2023c), learn new topics and languages (Jiang et al., 2024; Gupta et al., 2023), and even boost their general capabilities (Ibrahim et al., 2024). However, while many studies explore the mechanisms and properties of LLMs during pre-training from scratch (Biderman et al., 2023a; Xue et al., 2024), only a few research works focus on the behavior of LLMs during continual pre-training. To investigate the learning dynamic of LLM continual pre-training, we conduct experiments in the medical domain and closely monitor changes in model performance throughout the training process. Surprisingly, we find that the LLM’s performance on medical tasks drops at the early stage of training, despite a consistent improvement in perplexity on the medical corpus. However, as training progresses and more data are used, the task performance recovers and outperforms the original model. We further find that the initial performance drop and the following performance recovery phenomenon generally happen in LLMs across different scales and different training corpus.
Figure 1: The performance comparison between our model (Llama-3-physician) and other baselines involves reporting the ratio of each model’s task performance to the best performance of that task among all models.
To explain the LLM’s abnormal behavior during continual pre-training, we revisit the concept of the stability gap (De Lange et al., 2022; Caccia et al., 2021) from continual learning. The stability gap describes the phenomenon where the performance of old tasks initially drops and then recovers when learning a new task. A previous study (De Lange et al., 2022) explains this initial drop as the model having only a small stability gradient to maintain the performance of previous tasks compared to a large plasticity gradient for learning the new task at the beginning of continual learning. After the initial stage, the stability gradient will rise, leading to performance recovery. Following this framework, we hypothesize that the LLM’s performance on medical tasks depends on both medical domain knowledge and general instruction-following ability. Medical domain knowledge can be improved with the plasticity gradient, while the instruction-following ability is more related to the stability gradient. Therefore, the initial drop in medical task performance occurs due to an insufficient stability gradient to maintain the ability to follow instructions. To verify this hypothesis, we further study the model’s general domain task performance during continual pre-training on our medical-related corpus. The results show a similar v-shape curve as the medical domain tasks, where the task performance initially decreases and then starts to recover. Furthermore, from the perspective of weight updates, we find that the top layers’ weights, which contain high-level task knowledge, have comparatively small weight updates only at the initial stage. This also suggests that the LLM has only a small initial stability gradient to protect its instruction-following ability.
The stability gap causes inefficiency in continual pre-training as it delays the improvement in the LLM’s performance. To address this, we propose three efficient continual pre-training strategies:
  • 1.
    Instead of continually pre-training the LLM on a large corpus for one epoch, which induces a large plasticity gradient for a long period, we continually pre-train the LLM on a subset of the corpus with a proper size for multiple epochs.
  • 2.
    Select the subset with the highest-quality tokens to learn rich domain knowledge, leading to faster performance recovery and higher peak performance.
  • 3.
    Use a data mixture rate similar to the pre-training data, thus reducing the distribution shift and mitigating the knowledge forgetting of general instruction-following ability.
To verify our strategy, we first conduct experiments on the OpenLlama-3B model. We find that our strategies not only accelerate performance improvement by mitigating the stability gap but also improve the LLM’s peak performance. We also compare our strategies with other continual pre-training techniques and analyze the influence of important learning factors, such as learning rate, for our strategies in Sec. 5. Finally, we apply our strategies to both the continual pretraining and instruction tuning processes of the Llama-3-8B model (Meta, 2024), efficiently enhancing its performance on diverse medical tasks, outperforming other open-source LLM baselines, and achieving performance comparable to GPT-4 (See performance comparison in Figure 1).
2 Related work
Large language Models
such as GPT-4 (OpenAI, 2023), Gemini (Team, ), and Llama (Touvron et al., 2023a)), have billions of parameters and show strong performance on various basic natural language tasks (Qin et al., 2023), human examination (Hendrycks et al., 2020b; Zhong et al., 2023), and agent-related tasks (Guo et al., 2023; Liu et al., 2023; Zhou et al., 2023). Their success attracts researchers to analyze LLMs’ learning properties during the pre-training process (Kaplan et al., 2020; Biderman et al., 2023a; Zhang et al., 2024a).  Kaplan et al. (2020) finds the pre-training scaling rule for model size and dataset size and then Hoffmann et al. (2022) proposes the Chinchilla rule that claims the equal importance of the model size and the number of training tokens.  Sorscher et al. (2022) further claims that pruning low-quality data can improve the above neural scaling laws. However, high-quality training tokens are limited and may be run out soon (Villalobos et al., 2022). Thus, some researchers try to maximize the utilization of the existing corpus by training it for multiple epochs (Muennighoff et al., 2024; Xue et al., 2024). But they observe the performance degradation (Hernandez et al., 2022; Xue et al., 2023; Hoffmann et al., 2022) after training 4 epochs.
Continual pre-training
gradually becomes necessary for LLMs to expand their basic ability (Wu et al., 2022; Fu et al., 2024; Zhuang et al., 2024), avoid outdated information (Jiang et al., 2024), and become the domain expert (Huang et al., 2023; Yang et al., 2024a; Chen et al., 2023c; Nguyen et al., 2023; Wu et al., 2023; Yıldız et al., 2024; Xie et al., 2024a). The domain corpus for continual pre-training can be collected by n-gram models (Muennighoff et al., 2024), heuristic rules designed by human experts (Chen et al., 2023c; Zhang et al., 2024c) or automatically identified by a LLM (Zhang et al., 2024c). For the continual pre-training techniques.  Ke et al. (2023, 2022) focused on adding masks or adjusting the architecture of small Language models like RoBERT to protect the learned general knowledge. However, these techniques result in huge computational consumption for LLMs. Recent studies (Gupta et al., 2023) show that learning rate re-warming can improve LLMs’ downstream task performance.  Ibrahim et al. (2024) further claims that learning rate re-warming, re-decaying, and replay can make the continual pre-training performance match the performance of fully re-training when continually pre-training the English LLM on the German corpus. Other continual pre-training method studies focus on selecting useful tokens (Lin et al., 2024), expanding MOE architecture (Chen et al., 2023a), and knowledge distillation (Jin et al., 2021b).
Continual learning and the Stability Gap
Continual learning aims to design methods that can learn new knowledge without the catastrophic forgetting of previously learned knowledge (Kirkpatrick et al., 2017; Van de Ven et al., 2022). To mitigate the forgetting problem when learning a new task, replaying previous tasks’ data (Rolnick et al., 2019; Buzzega et al., 2020; Prabhu et al., 2020; Buzzega et al., 2021; Guo et al., 2022) becomes the main approach.  De Lange et al. (2022); Caccia et al. (2021) further find that, although they conduct the replay approach, the vision model still first loses its performance stability in previous classification tasks ( the performance drops abruptly) and then gradually recovers. They call it the stability gap phenomenon. Different from them, we focus on the continual pre-training of the LLM and observe that both the LLM’s domain task performance and general ability suffer from the stability gap.
3 Identifying the stability gap in continual pre-training
In this section, we first study the behavior of large language models (LLMs) during continual pre-training by measuring them at regular intervals. We observe that the performance on the target task initially drops and then rises during continual pre-training. To explain our observations, we introduce the concept of the stability gap to continual pre-training and verify our explanations with experiments.
3.1 Investigating the behavior of LLMs during continual pre-training
Experiment setup
In this study, we chose OpenLlama3B-v2 (Geng and Liu, 2023) as our default LLM and use the medical domain as our primary target domain. Following previous work (Chen et al., 2023b), we set the compute budget to 50 billion (50B) training tokens. To collect the continual pre-training corpus, we follow the simple and scalable methodology of Muennighoff et al. (2024). First, we train a KenLM model (Heafield, 2011) on a high-quality medical reference corpus. Then, we use the trained KenLM model to calculate the perplexity (PPL) of samples in the Refined-Web dataset (Penedo et al., 2023). Finally, we extract 50B tokens from the Refined-Web dataset with the lowest PPL to create the medical corpus. More details are provided in Appendix A.
Figure 2: (a) reports the models’ average medical performance during the medical continual pre-training process. (b) illustrates the models’ average medical perplexity (PPL) during the medical continual pre-training process. (c) shows the Pythia model’s average common-sense task performance when we continually pre-train it on the new Refined-Web datasets.
Observation (1): The medical task performance first drops and rises during continual pre-training.
Specifically, we follow Chen et al. (2023c) and measure the average accuracy performance over the MMLU-Medical-Genetics (Hendrycks et al., 2020a), MedQA (Jin et al., 2021a), PubMedQA (Jin et al., 2019), and MedMCQA (Pal et al., 2022) tasks (see task details in Appendix B). We report the average performance on medical tasks every 5 billion training tokens. From Figure 2(a), we observe that the domain task performance initially drops during the first 5 billion tokens and then gradually recovers and improves. Additionally, we consider the TinyLlama model (Zhang et al., 2024b), a 1.1B Llama model trained on 3 trillion tokens, and continually pre-train it on the medical corpus. From Figure 2(a), we observe that its performance on medical tasks also shows the same trend, despite being trained on so many tokens.
Observation (2): The perplexity of medical Wikipedia steadily declines during continual pre-training.
We further measure the average perplexity (PPL) of the models on the Wikipedia corpus about medical terms [^1] https://huggingface.co/datasets/gamino/wiki_medical_terms. From Figure 2(b), we observe that the PPL steadily drops. This indicates that the LLM has acquired medical domain knowledge at the initial continual pre-training and continues improving its medical domain knowledge throughout the entire continual pre-training process.
Observation (3): The general task performance also first drops and then rises during general continual pre-training.
Continual pretraining on another large corpus is an important approach to boost the pretrained LLM’s general task performance (Jiang et al., 2024; Gupta et al., 2023). We call it the general continual pretraining setting. We further find that it also exists a similar performance phenomenon. Specifically, we continually pre-train the Pythia-410m model (Biderman et al., 2023b) (initially pre-trained on the Pile (Gao et al., 2020) dataset) on the RefinedWeb dataset (Penedo et al., 2023) to boost its general ability. We measure its general ability using the average performance across 10 common-sense tasks and report the average performance of every 10 billion tokens. Training details are in Appendix A and task details are in Appendix B. From Figure 2(c), we observe that the LLM’s general task performance first drops significantly and then gradually rises.
Based on our observations, the initial drop followed by a rise in target task performance is a general phenomenon in the continual pre-training of LLMs of various sizes. This abnormal behavior can be explained with the concept of the stability gap (Lange et al., 2022) in the following section.
Figure 3: (a) shows the OpenLLaMa’s average common-sense task performance during medical continual pre-training. (b) illustrates the OpenLlama model’s relative parameter update during the medical continual pre-training process. We report the average weight relative update of weights in the top 5 layers and the bottom 5 layers. We also report the rate between the two average numbers.
3.2 Stability Gap: A conceptual explanation for the initial performance drop and then following recovery.
The stability gap
phenomenon has been observed as the model’s performance in previous tasks first drops and then rises when learning the new task and  Lange et al. (2022) explains it by disentangling the gradient $\mathcal{L}$ in $\alpha$-weighed plasticity and stability gradients: $\mathcal{L}=\alpha\mathcal{L}_{\textit{plasticity}}+(1-\alpha)\mathcal{L}_{\textit{stability}}$ where $\mathcal{L}_{\textit{plasticity}}$ aims to learn to the new task and $\mathcal{L}_{\textit{stability}}$ maintains the previous tasks’ performance. The performance of previous tasks initially drops because the plasticity gradient exceeds the stability gradient, causing neglect in maintaining prior task performance. Subsequently, the performance loss in previous tasks enhances the stability gradient, while adaptation to the new task reduces the plasticity gradient, leading to gradient balance and performance recovery.
Explanation of our observations
We assume that the performance on the new domain task relies on both the target domain knowledge and the instruction-following ability during continual pre-training. Inspired by the stability gap explanation framework, we decouple the continual pre-training gradient into the plasticity gradient for learning new domain knowledge and the stability gradient for maintaining instruction-following ability. Initially, the stability gradient is small compared to the plasticity gradient, causing a performance drop on the target domain because of the destructed instruction-following ability. Subsequently, the stability gradient increases to recover the instruction-following ability, while the plasticity gradient already helps the LLM learn domain knowledge, observing performance improvement on the target domain.
Empirical verification for our explanation
We verify our explanation by (1) measuring general task performance during continual pre-training on the medical-related corpus. As shown in Figure 3(a), the general task performance follows a similar V-shape curve, indicating the recovery of general instruction-following ability after the initial drop. We also verify our explanation at the weight level by (2) measuring the relative weight update of each weight $w$ as $\frac{w_{\textit{t}}-w_{0}}{w_{0}}$, where $w_{\textit{t}}$ is the weight value during continual pre-training and $w_{0}$ is the original weight value. A high relative weight update indicates a large gradient for updating the weight. Figure 3(b) shows that the bottom layers’ weights initially have a higher relative weight update than the top layers (rate > 1.35). Previous studies indicate that bottom layers learn the syntax and low-level semantics (Devlin et al., 2019; Hewitt and Manning, 2019; Ling et al., 2023), while top layers contain high-level semantics and task-specific knowledge (Yang et al., 2024b; Chen et al., 2024). This suggests that the top layers’ weights indeed lack sufficient stability gradient to maintain instruction-following ability initially. The performance then recovers as the relative weight updates (stability gradient) increase in the top layers and domain knowledge is learned, as indicated by the continuous drop in medical perplexity.
4 Efficient continual pretraining strategies for mitigating the stability gap
In this section, we propose three efficient continual pre-training strategies for overcoming the above stability gap problem. The training process and details follow those in the above section.
Figure 4: (a) reports the average medical performance during the medical continual pre-training process. The baseline is pre-training the OpenLlama-3B model with 50b medical tokens with one epoch. ’5b Random’ is pre-training the LLM with 5b tokens randomly selected from the 50b medical tokens for 5 epochs. ’5b HQ’ is pre-training the LLM with the highest quality (HQ) 5b tokens of the 50b medical tokens for 5 epochs. (b) shows the average medical performance across 5 epochs. (c) illustrates the average commonsense task performance across 5 epochs.
Strategy I: Continually pre-train the LLM on a corpus subset across multiple epochs rather than the entire large corpus for a single epoch.
The domain task performance depends on the general instruction-following ability, which is related to the rise of the stability gradient. The usual continual pre-training approach collects as many samples as possible and trains them one time. However, this approach means the LLM maintains a high plasticity gradient for learning new samples in every batch, causing a stability gap and a slow rise in the stability gradient. To mitigate this, we propose randomly selecting a subset with a proper size from the entire corpus and pre-training the LLM on this subset across multiple epochs. This method reduces the need for a high plasticity gradient after the first epoch and accelerates the rise of the stability gradient (performance recovery). In Figure 4(a), we observe that this strategy leads to faster performance recovery. The LLM achieves peak performance at the fourth epoch, consistent with previous studies (Xue et al., 2024).
Strategy II: Continually pre-train the LLM on the corpus subset with the highest quality.
The performance of domain tasks also depends on the learned domain knowledge. Therefore, collecting a subset with the highest quality should further enhance performance. To verify this, we used the trained KenLM from Sec. 3.1 to calculate the perplexity (ppl) of each sample in the entire medical corpus. A lower perplexity indicates that the sample is closer to the distribution of the medical reference corpus. We then continually pre-trained the OpenLlama-3B model on the subset with the lowest perplexity (i.e., the highest quality) for multiple epochs. From Figure 4 (a), we observe that the high-quality subset indeed enables the LLM to recover performance faster and stronger in the medical domain. Further analysis of the pre-training subset size is presented in Sec. 5.2.
Strategy III: Use a data mixture rate similar to the pre-training data.
The pre-training data mixture rate is a vital factor for the pre-training performance of large language models (LLMs) (Xie et al., 2024c; Shen et al., 2023). Therefore, we propose a third strategy that follows the pre-training data’s mixture rate to construct the continual pretraining training subset, aiming to reduce the distribution gap and stabilize the instruction-following ability of the LLM during continual pre-training. Specifically, for the OpenLlama model, we follow the Llama mixture rate (Touvron et al., 2023a) to collect 5 billion tokens initially. We then replace the CC and C4 data (82% of the 5 billion tokens) with medical tokens sampled from the highest quality 5 billion medical tokens (HQ-5b). There are two ways to sample these medical tokens. The first method randomly samples the medical tokens once to construct a fixed training corpus. We call this “rate-fixed-data-fixed”. The second method randomly samples the medical tokens from the HQ-5b tokens for each epoch. We call this “rate-fixed-data-dynamic”.
From Figure 4(b), we observe that both methods improve the performance of the first epoch by overcoming the stability gap. The second method achieves a higher peak performance as it offers a better trade-off between recovering the instruction-following ability through data replay and learning domain knowledge from a large number of medical tokens. Additionally, our strategies further improve the average performance on general commonsense tasks, as shown in Figure 4(c), and reduce the medical perplexity and the rate of relative weight update, as detailed in Appendix C. We also investigate the effectiveness of our three strategies in the general continual pre-training setting in Appendix D.
5 Evaluation
In this section, we first compare the effectiveness of our strategies with other continual pre-training techniques. Next, we investigate the impact of important learning factors, such as the learning rate, on our strategies. Finally, we deploy our strategies into the newest Llama-3-8b model, which achieves the strongest fine-tuned performance among open-source baselines.
5.1 Comparison with other continual pre-training techniques
Baselines and evaluation tasks
We consider the following baselines for comparison: (1) Continually pre-training the OpenLLaMa-3B LLM with 50 billion collected medical tokens for one epoch ("the full token baseline"). (2) Re-warming and re-decaying the learning rate of (1) based on the paper by (Ibrahim et al., 2024). (3) Replay baselines: Following (Chen et al., 2023b), we replace 10% of the tokens in (1) with tokens randomly sampled from the OpenLLaMa-3B’s pre-training dataset (the RefinedWeb dataset). We also consider a replay baseline that replaces 10% of medical tokens in the experiment using strategies I and II with randomly selected tokens from the RefinedWeb dataset. This baseline does not consider the data mixture rate. (4) Parameter protection baselines: Following (Harun and Kanan, 2023), we freeze the top 5 layers’ weights during the continual pre-training process of (1) to protect the high-level instruction-following ability and mitigate the stability gap. We also consider another baseline that freezes the bottom 5 layers’ weights for comparison. We follow (Chen et al., 2023b) and consider the tasks of PubMedQA, MedMCQA, and MedQA-4-Option. For the MMLU benchmark (Hendrycks et al., 2020a), we consider the average performance of its medical topics, including medical genetics, anatomy, clinical knowledge, professional medicine, and college medicine. We use the lm-evaluation-harness framework (Gao et al., 2023) to measure the baselines’ zero-shot performance.
Method Training tokens number MMLU-Med-Avg PubMedQA MedMCQA MedQA-4-Option Avg
OpenLLaMa-3B - 25.6 68.4 25.4 25.4 36.2
Full token baseline 50B 26.1 70.4 26.1 27.1 37.4
Re-warming and re-decaying 50B 26.5 70.3 27.1 27.1 37.7
Replay 10% data 50B 26.3 69.2 27.6 26.9 37.5
Replay 10% data with our strategies 20B 29.3 71.0 30.4 27.6 38.5
Freezing top 5 layers 50B 26.2 69.9 27.1 27.3 37.6
Freezing bottom 5 layers 50B 26.0 69.1 25.4 25.7 36.5
Our strategies 20B 30.0 71.2 34.0 27.8 40.7
Table 1: Zero-shot accuracy across various medical benchmarks.
Results
From Table 1, we find that (1) our strategies improve the base model’s average medical task performance significantly (4.5%) with only 20 billion training tokens. This demonstrates the effectiveness and efficiency of our strategies for continual pre-training. (2) Other techniques can also improve continual pre-training performance, except for the baseline ’Freezing bottom 5 layers,’ which hinders the learning of medical domain knowledge. (3) Following the pre-training data mixture rate to replay pre-training data is more effective than randomly sampling pre-training data for replay. This is because it further reduces the distribution shift between the pre-training corpus and the continual pre-training corpus, thereby helping to recover the LLM’s general instruction-following ability.
5.2 Analysis
Figure 5: (a) reports the performance of TinyLlama-1.1B across multiple epochs. All these experiments use our strategies with different pre-training learning rates. (b) reports the performance of OpenLlama-3B across multiple epochs. All of the experiments in (a) and (b) use our strategies with different pre-training learning rates. (c) reports the performance of OpenLlama-3B across multiple epochs with different training subset sizes $S$. To collect the pre-training corpus with different sizes, we first rank all samples of the 50 billion medical tokens based on the perplexity calculated by the trained KenLM (see Sec. 3.1). Then, we select the first $S$ billion tokens with the lowest perplexity. For all experiments here, we report the average task performance of PubMedQA, MedMCQA, MMLU-medical-genetics, and MedQA-4-Option tasks.
Impact of learning rate
The pre-training learning rate is a crucial factor for updating LLMs during continual pre-training. To investigate its impact on our strategies, we conduct continual pre-training experiments with different learning rates. From Figure 5(a) and (b), we find that the optimal learning rate varies with the LLM scale: a small LLM (e.g., TinyLlama-1.1B) requires a higher learning rate (e.g., 3e-4), whereas larger LLMs (e.g., OpenLlama-3B) benefit from a lower learning rate (e.g., 3e-5). If the learning rate is too low (e.g., 3e-5 for TinyLlama-1.1B), the LLM cannot learn domain knowledge effectively to boost performance. Conversely, if the learning rate is too high (e.g., 3e-4 for OpenLlama-3B), performance declines as the large learning rate leads to a significant plasticity gradient, causing the LLM to lose its general instruction-following ability for completing tasks. Based on our analysis experiments, we set the pre-training learning rate at 3e-4 for TinyLlama and 3e-5 for OpenLlama-3B’s experiments.
Impact of the training subset size
The size of the training subset is another important factor in our strategies. To determine the optimal training subset size, we conduct pre-training experiments on Llama-3b using various training subset sizes. From Figure 5(c), we observe that a smaller high-quality subset yields better initial performance and mitigates the stability gap (e.g., 1 billion tokens), but it also causes the performance to drop quickly in later epochs due to overfitting. A larger subset (e.g., 10 billion tokens) results in a stability gap and slower performance recovery, as the LLM needs to maintain a high plasticity gradient to learn a large number of new samples. Based on our experiments, we select a subset with 5 billion high-quality tokens, as it mitigates the stability gap, achieves the best peak performance, and is computationally effective.
5.3 Deploying our strategies into the Llama-3 Model
Continual pre-training
We continually pre-train the Llama3-8B-base model using our three strategies with the high-quality 5 billion medical tokens constructed in Sec. 4 for 4 epochs. The training details are in Appendix E. After the continual pre-training process, we find that the average medical performance drops slightly, likely due to the unknown data mixture rate of Llama-3 and the lack of access to its high-quality pre-training corpus for performance recovery. However, the medical perplexity is significantly lower than that of the Llama3-8B-base model.
Task-specific fine-tuning
To evaluate LLMs’ performance in the supervised learning setting, we follow (Chen et al., 2023b) and individually conduct task-specific finetuning on both the base models and the continually pre-trained models using each benchmark’s training set. Since MMLU (Hendrycks et al., 2020a) does not have a training set, we follow (Chen et al., 2023b) and primarily consider the MMLU-Medical-Genetics benchmark, evaluating the model finetuned on MedMCQA. We put task details in Appendix B and training details in Appendix E.
Baselines
For task-specific fine-tuning, we consider three kinds of baselines here: (1) Task-specific finetuning of the base model of open-source LLMs. This includes models such as Llama-2-70B, Llama-3-8B, and Llama3-Aloe-8B-Alpha (Gururajan et al., 2024). We copy their results from their respective papers (Gururajan et al., 2024) or the Meditron paper (Chen et al., 2023b) except for the Llama-3-8B, which we finetuned using the same process as our strategies. (2) Task-specific finetuning of continually pre-trained LLMs like meditron (Chen et al., 2023b), BioMistral SLERP 7B (Labrak et al., 2024), Llama-3-8B-full. These LLMs have been continually pre-trained with a medical corpus. We copy their results from their papers, except for Llama-3-8B-full, for which we continually pre-train the Llama-3-8B with 50B medical tokens collected in Section 3.1, and then finetune it using the same process as our strategies. (3) Closed-source LLMs. This includes models like ChatGPT and GPT-4 (OpenAI, 2023). The results are measured using the Microsoft Azure OpenAI API service (Shi et al., 2024).
Model MMLU-Medical PubMedQA MedMCQA MedQA-4-Option Avg
Llama-2-7B (Touvron et al., 2023b) 56.3 61.8 54.4 49.6 53.2
BioMistral SLERP 7B (Labrak et al., 2024) 60.5 75.2 44.2 47.3 56.8
MEDITron-7B (Chen et al., 2023b) 55.6 74.4 59.2 52.0 57.5
Llama3-Aloe-8B-Alpha (Gururajan et al., 2024) 72.7 77.2 59.0 62.3 67.8
Llama-2-70B 74.7 78.0 62.7 61.3 67.2
MEDITron-70B 73.6 80.0 65.1 65.4 69.0
GPT-3.5-turbo-finetuned (Shi et al., 2024) 70.5 71.4 61.8 63.3 66.7
Llama-3-8B Fine-tuned (ours) 82.3 75.8 60.0 61.1 69.8
Llama-3-8B Full (ours) 82.0 78.6 61.8 60.8 70.8
Llama-3-Physician-8B (ours) 85.0 79.1 81.4 61.5 76.7
Table 2: Accuracy comparison across various medical benchmarks in the task-specific fine-tuning setting. Llama-3-8B Fine-tuned is directly fine-tuned on these tasks. For ’Llama-3-8B Full’, we first continually pre-trained the Llama with 50B medical tokens and then finetuned the pretrained model on these tasks. For Llama-3-Physician-8B, we first continually pre-trained the Llama with with our strategies and then finetuned the pretrained model on these tasks.
Results
We use the lm-eval-harness (Gao et al., 2023) to evaluate our model (Llama-3-Physician) and related baselines’ performance. No demonstration examples are used. From Table 2, we find that our model outperforms other baselines with similar model scales on the four evaluation benchmarks by a clear margin. This is due to the following reasons: (1) we use the newest and strongest open-source Llama-3 model rather than older Llama-2 or Mistral-7B, (2) we continually pre-train the base model with high-quality medical tokens (compared to ’Llama-3-8B fine-tuned and Llama-3-8B instruct’), and (3) our strategies further boost the gains from continual pre-training markedly (compared to ’Llama-3-8B Full’). Our model also outperforms many larger LLMs (70B) on average, meaning that users can obtain higher-quality medical services with a faster inference rate and less memory consumption.
5.4 Deploying our strategies into the instruction tuning process
Instruction-tuning is an important approach to boost the LLM’s performance among multiple tasks. We follow Xie et al. (2024b) and consider the instruction-tuning setting that tunes the continual pretrained Llama-3-8B model (see the above section) with a combination of question-answering tasks like PubMedQA (Jin et al., 2019), classification tasks like HOC (Baker et al., 2016), relation extract tasks like DDI2013 (Segura-Bedmar et al., 2013), inference tasks like BioNLI (Bastan et al., 2022), and summarization tasks like MIMIC-CXR (Johnson et al., 2019). The specific dataset details are in Appendix B. Unlike the above task-specific fine-tuning, we only tune one LLM here and use the instruction-tuned LLM to test all benchmarks. We tune the Llama model for 3 epochs with a learning rate of 3e-5. More training details are in Appendix E.
Figure 6: We consider the ’full instruction data’ experiment as fine-tuning the model with all instruction data for 3 epochs. For the ’$n$data’ experiments, we first uniformly sampled the highest quality instructions from each instruction dataset based on scores provided by the Deita data selector. We then mixed the sampled data with the general instructions from the Airoboros-3.2 dataset. The total training tokens are equal to $n$of the full instruction data. We set $n$ to 25, 50, and 75 here. (a) shows the experiments’ average medical question-answering task performance during instruction tuning. (b) illustrates the experiments’ performance for other medical tasks. For BioNLI, DDI 2023, and HOC tasks, we report macro-F1 as the score. For MIMIC-CXR summarization tasks, we report Rouge-L as the score.
Deployment
In the instruction tuning process, our first strategy is common as the medical instruction tuning process usually involves multi-epochs training (Zhang et al., 2023a; Xie et al., 2024b; Han et al., 2023). For the second strategy, we consider Deita (Liu et al., 2024), a simple automatic instruction data selector, to select high-quality medical instruction data. This selector uses the LLM to give quality scores for instructions and considers the diversity of instruction data by sampling data from different clustering. For the last strategy, we consider high-quality general instruction datasets like Airoboros-3.2 Durbin (2024) to mitigate the forgetting in general instruction following ability.
Observations
From Figure 6, we first observe that the average performance of medical question-answering tasks initially drops slightly (in the first epoch) and then gradually rises, which is similar to the phenomenon observed in the continual pre-training process. Additionally, we observe that our strategies can mitigate the initial performance drop and achieve higher peak performance during the instruction tuning process, thereby extending the application of our strategies. Figure 6 also shows that we only need computation equivalent to 25% of the original instruction data (consisting of high-quality medical instruction data and general instruction data) to achieve the best performance among diverse tasks. This reduces computational consumption and improves the efficiency of the instruction tuning process. We call the tuned model in the experiment ’25% instruction data’ as ’Llama-3-physician-8B instruct’. In the following paragraphs, we will compare it with other baselines.
Baselines
For instruction-tuning, we consider instruction-tuned models like Mistral-7B-instruct (Jiang et al., 2023), Zephyr-7B-$\beta$-instruct (Tunstall et al., 2023), PMC-Llama-7B (Wu et al., 2023), BioMedGPT-LM 7B (Zhang et al., 2023a), Medalpaca-13B (Han et al., 2023), AlpaCare-13B (Zhang et al., 2023b), Me-LLaMA-13B chat(Xie et al., 2024b), Llama-3-8B instruct (Meta, 2024), and JSL-Med-Sft-Llama-3-8B (johnsnowlabs, 2024). These LLMs are tuned with general instructions or medical task instructions.
Model MMLU-Medical PubMedQA MedMCQA MedQA-4-Option Avg
Mistral-7B-instruct (Jiang et al., 2023) 55.8 17.8 40.2 41.1 37.5
Zephyr-7B-instruct-$\beta$ (Tunstall et al., 2023) 63.3 46.0 43.0 48.5 48.7
PMC-Llama-7B (Wu et al., 2023) 59.7 59.2 57.6 49.2 53.6
Medalpaca-13B (Han et al., 2023) 55.2 50.4 21.2 20.2 36.7
AlpaCare-13B (Zhang et al., 2023b) 60.2 53.8 38.5 30.4 45.7
BioMedGPT-LM 7B (Zhang et al., 2023a) 52.0 58.6 34.9 39.3 46.2
Me-Llama-13B (Xie et al., 2024b) - 70.0 44.9 42.7 -
Llama-3-8B instruct 82.0 74.6 57.1 60.3 68.5
JSL-Med-Sft-Llama-3-8B (johnsnowlabs, 2024) 83.0 75.4 57.5 59.7 68.9
GPT-3.5-turbo-1106 74.0 72.6 34.9 39.3 60.6
GPT-4 (OpenAI, 2023) 85.5 69.2 69.5 83.9 77.0
Llama-3-physician-8B instruct (ours) 80.0 76.0 80.2 60.3 74.1
Table 3: Accuracy comparison for question-answering tasks in the instruction-tuning setting.
Task type Classification Relation extraction Natural Language Inference Summarization
Datasets HOC DDI-2013 BioNLI MIMIC-CXR
Mistral-7B-instruct (Jiang et al., 2023) 35.8 14.1 16.7 12.5
Zephyr-7B-instruct-$\beta$ (Tunstall et al., 2023) 26.1 19.4 19.9 10.5
PMC-Llama-7B (Wu et al., 2023) 18.4 14.7 15.9 13.9
Medalpaca-13B (Han et al., 2023) 24.6 5.8 16.4 1.0
AlpaCare-13B (Zhang et al., 2023b) 26.7 11.0 17.0 13.4
BioMedGPT-LM 7B (Zhang et al., 2023a) 23.4 15.5 17.9 6.2
Me-Llama-13B (Xie et al., 2024b) 33.5 21.4 19.5 40.0
JSL-Med-Sft-Llama-3-8B (johnsnowlabs, 2024) 25.6 19.7 16.6 13.8
Llama-3-8B instruct 31.0 15.1 18.8 10.3
GPT-3.5-turbo-1106 54.5 21.6 31.7 13.5
GPT-4 (OpenAI, 2023) 60.2 29.2 57.8 15.2
Llama-3-physician-8B instruct (ours) 78.9 33.6 76.2 37.7
Table 4: Performance comparison for general medical tasks in the instruction-tuning setting. For BioNLI, DDI 2023, and HOC tasks, we report macro-F1. For MIMIC-CXR summarization tasks, we report Rouge-L as the result.
Results
We download the baselines’ official models/deploy their APIs and then test their task performance using lm-eval-harnesses and Me-Llama’s evaluation frameworks. If the paper does not release its model, we copy the results from the original paper (e.g., Me-Llama). From Table 3, we find that our model outperforms other open-source baselines in question-answering tasks by a clear margin. Additionally, our model’s average performance is close to that of GPT-4. Furthermore, in Table 4, we observe that our model significantly outperforms GPT-4 in medical classification, relation extraction, natural language inference, and summarization tasks. This demonstrates the significant advantage of our model in processing diverse medical tasks. Finally, compared to these closed-source LLMs and larger open-source LLMs, our 8B model has the potential advantage of being deployed on users’ local devices, reducing the risk of leaking personal healthcare information.
6 Conclusion
Our paper explores the behavior of LLMs when continually pre-training them on a new domain’s corpus and observes the stability gap, a phenomenon marked by a significant initial performance drop followed by a slow recovery. We explain it from the view of plasticity and stability gradients and then propose three strategies that effectively improve the LLM’s domain performance and reduce computational costs by overcoming the stability gap. Furthermore, we deploy our strategies on the newest Llama-3-8B model, which achieves the strongest performance among open-source baselines of similar model scales and outperforms the closed-source GPT-3.5 model.
Limitations and Potential impacts
Ideally, knowing the pre-training data mixture could maximize the outcome of our method, but most strong open-source LLMs didn’t provide their training data mixture. Our Llama-3-8B experiment shows we can still improve significantly in this scenario. Due to limitations in computing resources, we plan to verify our conclusions and strategies on larger LLMs in the future. Our strategies are designed to address the machine learning problem of the stability gap, and we do not see any potential risks. The datasets and base models used in this paper will be open-sourced. Although we do not consider our model to be ready for real-world medical use in its current form, we are releasing it to the research community to promote work on large language models for the medical domain and the safety of language models in medical applications.
Appendix A The Details of Pre-training
For OpenLLaMa-3B, TinyLLaMa-1B, and Pythia-410m, we download them from their official website. For OpenLLaMa-3B and TinyLLaMa-1B LLMs, we continually pre-train them with the 50 billion medical tokens we constructed in Sec. 4 for one epoch. For the Pythia-410m LLM, we continually pre-train it with the 100 billion tokens randomly sampled from the 2021-2022 subset of the RefinedWeb dataset. We consider this subset as the Pile dataset only contains data before the year 2021 and then the tokens sampled from the 2021-2022 subset are unseen for the Pythia-410m model. The pre-training code is based on the transformers. The task is to predict the next token with a context size of 2048. The training is executed using 192 V100 GPUs. We employ the AdamW optimizer with $\beta_{1}=0.9,\beta_{2}=0.95$, a weight decay of 0.01, and a learning rate of 3e-4. We use a cosine learning rate scheduler with a 0.1 warmup ratio for gradual adaptation to training complexity and bf16 precision for computational efficiency. Gradient accumulation is set to 4 steps, and each training batch contains about 340 million tokens. We also add support for FlashAttention-2  [Dao, 2023] for more efficient inference and long-context decoding.
Appendix B Task Information
For the medical evaluation, we follow Chen et al. [2023b] and mainly consider the following four tasks:
MedMCQA [Pal et al., 2022] is a large-scale and comprehensive dataset for multichoice (four-option) medical question answering. It is derived from real-world medical entrance exam questions (Indian AIIMS and NEET-PG) and consists of over 194,000 high-quality medical questions. These questions cover 2,400 healthcare topics and 21 medical subjects, exhibiting a wide range of topical diversity. The average token length is 12.77.
MedQA [Jin et al., 2021a]is a multichoice question-answering dataset collected from the professional medical board exam, the United States Medical License Exams (USMLE). It comprises 12,723 questions sourced from a comprehensive collection of 18 English medical textbooks that have been extensively utilized by medical students and USMLE candidates. Questions in MedQA cover a wide range of topics in clinical medicine, necessitating responses with professional expertise and complex multi-hop reasoning across multiple pieces of evidence. The average question and option length is 116.6 and 3.5, respectively.
MMLU [Hendrycks et al., 2020b] is a comprehensive multi-task language understanding test dataset that encompasses 57 tasks across various domains such as mathematics, history, computer science, law, etc. In our experiments, we specifically focus on a subset of medical reasoning-related tasks including clinical knowledge, college medicine, medical genetics, and professional medicine.
PubMedQA [Jin et al., 2019] is a biomedical question and answering dataset derived from PubMed abstracts. It contains 1k expert annotated multi-choice question-and-answer samples based on 211.3k PubMed articles. The task of PubMedQA is to provide answers to research questions with yes/no/maybe responses based on the corresponding abstracts. The average question and context length is 14.4 and 238.9, respectively.
HOC [Baker et al., 2016] is a classification task to decide the Hallmarks of Cancer (HOC) taxonomy of the article based on its abstract. The input is an abstract text. There are 10 topics you will need to decide whether the article is related to. Topics: sustaining proliferative signaling, evading growth suppressors, resisting cell death, enabling replicative immortality, inducing angiogenesis, activating invasion and metastasis, genomic instability and mutation, tumor promoting inflammation, and cellular energetics, and avoiding immune destruction.
[…]
BioNLI [Bastan et al., 2022] is a task to classify the relationship between the given medical premise and hypothesis into one of the following labels: entailment, contradiction, or neutral. This dataset contains abstracts from biomedical literature and mechanistic premises generated with nine different strategies.
MIMIC-CXR [Johnson et al., 2019] is a generation task that derives the impression from findings in the radiology report.
The dataset statistics are in Table 5
Table 5: Dataset statistics
Dataset # Train # Test Source
MedMCQA [Pal et al., 2022] 182,822 4183 Exam
MedQA [Jin et al., 2021a] 10178 1273 Exam
MMLU [Hendrycks et al., 2020b] - 163 Exam
PubMedQA [Jin et al., 2019] 211,269 500 Literature
HOC [Baker et al., 2016] 1108 315 Literature
DDI 2023 [Segura-Bedmar et al., 2013] 1108 315 Literature
BioNLI [Bastan et al., 2022] 5544 6308 Literature
MIMIC-CXR [Bastan et al., 2022] 122,014 1606 Literature
For the evaluation of general task ability, we consider the following 10 commonsense tasks:
ARC-Challenge and ARC-Easy ARC [Clark et al., 2018] is a multiple-choice question-answering dataset, containing questions from science exams from grade 3 to grade 9. The dataset is split into two partitions: Easy and Challenge, where the latter partition contains the more difficult questions that require reasoning. Most of the questions have 4 answer choices.
BoolQ [Clark et al., 2019] is a question-answering dataset for yes/no questions containing 15942 examples. These questions are naturally occurring —they are generated in unprompted and unconstrained settings. Each example is a triplet of (question, passage, answer), with the title of the page as optional additional context. The text-pair classification setup is similar to existing natural language inference tasks.
COPA [Roemmele et al., 2011] consists of 1000 questions, split equally into development and test sets of 500 questions each. Each question is composed of a premise and two alternatives, where the task is to select the alternative that more plausibly has a causal relation with the premise.
HellaSWAG [Zellers et al., 2019] is a dataset for studying grounded commonsense inference. It consists of 70k multiple choice questions about grounded situations: each question comes from one of two domains – activitynet or wikihow – with four answer choices about what might happen next in the scene. The correct answer is the (real) sentence for the next event; the three incorrect answers are adversarially generated and human-verified, so as to fool machines but not humans.
OpenBookQA [Mihaylov et al., 2018] is a new kind of question-answering dataset modeled after open-book exams for assessing human understanding of a subject. It consists of 5,957 multiple-choice elementary-level science questions (4,957 train, 500 dev, 500 test), which probe the understanding of a small “book” of 1,326 core science facts and the application of these facts to novel situations.
PIQA [Bisk et al., 2020] dataset introduces the task of physical commonsense reasoning and a corresponding benchmark dataset Physical Interaction: Question Answering or PIQA. Physical commonsense knowledge is a major challenge on the road to true AI-completeness, including robots that interact with the world and understand natural language. PIQA focuses on everyday situations with a preference for atypical solutions.
Race [Lai et al., 2017] is a large-scale reading comprehension dataset with more than 28,000 passages and nearly 100,000 questions. The dataset is collected from English examinations in China, which are designed for middle school and high school students. The dataset can serve as the training and test sets for machine comprehension.
SciQ [Welbl et al., 2017] dataset contains 13,679 crowdsourced science exam questions about Physics, Chemistry and Biology, among others. The questions are in multiple-choice format with 4 answer options each. For the majority of the questions, an additional paragraph with supporting evidence for the correct answer is provided.
WinoGrande [Sakaguchi et al., 2021] is a new collection of 44k problems, inspired by the Winograd Schema Challenge (Levesque, Davis, and Morgenstern 2011), but adjusted to improve the scale and robustness against the dataset-specific bias. Formulated as a fill-in-a-blank task with binary options, the goal is to choose the right option for a given sentence which requires commonsense reasoning.
We use the lm-eval-harness [Gao et al., 2023] to evaluate the LLM on these tasks’ test set and report the zero-shot performance.
Appendix C The Perplexity and relative parameter update rate of the LLM using our strategies
Figure 7: (a) reports the average medical perplexity of the OpenLLaMa-3B using our strategies. ’5b HQ’ means the LLM using our strategies I and II. ’5b rate-fixed-data-dynamic’ means the LLM using our three strategies. ’Baseline’ is the average medical perplexity of the OpenLLaMa-3B model that has been continually pre-trained with 50 billion medical tokens. (b) shows the rate between the bottom 5 layers’ average relative parameter and the top 5 layers’ average relative parameter update of the OpenLLaMa-3B using our strategies. ’Baseline’ is the rate of the OpenLLaMa-3B model during the continual pre-training with 50 billion medical tokens.
From Figure 7(a), we observe that the LLM using our strategies gradually decreases its average medical perplexity, indicating that the LLM is acquiring rich medical knowledge. Its average medical perplexity at the fourth epoch is even lower than that of the OpenLLaMa-3B model, which has been continually pre-trained with 50 billion medical tokens. From Figure 7(b), we also find that the ratio between the average relative parameter updates of the bottom 5 layers and the top 5 layers of the OpenLLaMa-3B model using our strategies is closer to 1. This suggests that the plasticity gradient and the stability gradient are more balanced when employing our strategies.
Appendix D Deploying our strategies into the general continual pre-training setting
Figure 8: We report the average performance of the 10 commonsense and reading compression task here. The Model is Pythia-410m.
Continually pre-training one LLM on another large corpus is an approach to boost its general ability [Gupta et al., 2023]. We consider the scenario of continually pre-training the Pythia-410m model on the RefinedWeb dataset. The Pythia-410m model has been pre-trained on the Pile dataset. In this context, we use the average performance of 10 commonsense and reading comprehension tasks, as detailed in Appendix B, to measure the LLM’s general task performance. To test the effectiveness of strategy I in the general continual pre-training setting, we conduct multi-epoch experiments with different training subset sizes. The tokens in each training subset are randomly sampled from the RefinedWeb dataset and the computational consumption of each experiment can not be beyond the compute budget (100 billion tokens). From Figure 8, we find that strategy I indeed helps the Pythia-410m model to mitigate the stability gap and achieve better peak performance. We also find the best performance among our experiments is achieved when pre-training the LLM with 11 billion tokens for 7 epochs. However, we can not find a good quality filter for the second strategy. We have tried to train a KenLM on WikiText as the quality filter for measuring the sample’s quality in improving LLMs’ general ability. But it does not work. From Figure 8, we find that strategies I and III can help the LLM to overcome the stability gap and achieve higher performance.
Appendix E The Training Details of Deploying our Strategies into the Llama-3 Model
Pre-training details: The pre-training task is to predict the next token with a context size of 8192. The training is executed using 16 H100 80GB GPUs. We employ the AdamW optimizer with $\beta_{1}=0.9,\beta_{2}=0.95$, a weight decay of 0.01, and a learning rate of 3e-5. We use a cosine learning rate scheduler with a 0.1 warmup ratio for gradual adaptation to training complexity and bf16 precision for computational efficiency. Gradient accumulation is set to 12 steps, and each training batch contains about 340 million tokens. We also add support for FlashAttention-2  [Dao, 2023] for more efficient inference and long-context decoding.
Task-specific finetuning details: We employ the AdamW optimizer with a weight decay of 0.01 and a learning rate of 3e-5. We use a cosine learning rate schedule with a 10% warmup ratio, decaying the final learning rate to 10% of the peak learning rate. We fine-tune the LLMs for 3 epochs.
Instructions-tuning details: We consider the combination of the training set of MedMCQA [Pal et al., 2022], MedQA [Jin et al., 2021a], PubMedQA [Jin et al., 2019], HOC [Baker et al., 2016], DDI2013 [Segura-Bedmar et al., 2013], BioNLI [Bastan et al., 2022], and MIMIC-CXR [Johnson et al., 2019] tasks . To avoid potential data contamination, for each test sample of MedQA [Jin et al., 2021a], PubMedQA [Jin et al., 2019], and MedMCQA [Pal et al., 2022] tasks, we delete the training samples that contain its option. For the training samples of theMedQA [Jin et al., 2021a],PubMedQA [Jin et al., 2019], and MedMCQA [Pal et al., 2022] tasks, we use the instruction template from the Meditron paper [Chen et al., 2023b]. For the other datasets’ training samples, we use their original instructions. We employ the AdamW optimizer with a weight decay of 0.01 and a learning rate of 3e-5. We use a cosine learning rate schedule with a 10% warmup ratio, decaying the final learning rate to 10% of the peak learning rate. We fine-tune the LLMs for 3 epochs. The global batch size is 96 and max sequence length is 1024.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and precede the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our abstract and introduction clearly state the claims made, including the contributions made in the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Yes, we discuss our limitations after the conclusion section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
[…]
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We list all experimental details in the evaluation section and appendixes. We will release our models in public (e.g., HuggingFace).
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The pre-training data is too large to put it into the material. The fine-tuning data is open-source. We will release the code and relevant data to reproduce our results soon.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperweights, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We write all the training and test details in the evaluation section and appendixes.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
[…]
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See the evaluation section and appendixes.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We conform with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Yes, please read the paragraph after the conclusion section.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pre-trained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: Yes, we describe it in the paragraph after the conclusion section.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Yes, we properly credit them.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Yes, we introduce the details of our new models in the main paper.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.